OpenStudy (anonymous):

solve Γ 5/2 Γ 3 --------- Γ 11/2 dont give just answer plz solve it

OpenStudy (anonymous):

\[\lceil n =(n-1)!\]

OpenStudy (anonymous):

\[\lceil (1/2)=\sqrt{\pi}\]